# PointNSP: Autoregressive 3D Point Cloud Generation with Next-Scale Level-of-Detail Prediction

### v2

@@ 4 Experiments
Table 1: The Performance (1-NNA) is evaluated based on single-class generation. The second block specifies the types of generative models used in each study. The best performance is highlighted in bold, while the second-best performance is underlined. Performance is reported on two dataset splits: the top corresponds to the random split, and the bottom corresponds to the LION split.
@@
Table 2: Training time (in GPU hours, averaged over three categories), sampling time (in seconds, averaged over $10$ samples), and model size (in millions of parameters). Ranked by MeanCD and MeanEMD.
  Quality Rank Model
Training Time
(GPU hours) $\downarrow$
Sampling Time
(s) $\downarrow$
Param (M) $\downarrow$
8 PointGrow 156 5.80 25
7 CanonicalVAE 142 5.45 30
6 PointGPT 185 5.32 46
5 PVD 142 29.9 45
4 LION 550 31.2 60
3 TIGER 164 23.6 55
[…]
1 PointNSP-m (ours) 178 3.59 32
4.1 Experimental Setup
Datasets.
In line with prior studies, we use ShapeNetv2, pre-processed by PointFlow [PointFlow], as our primary dataset. Following previous experimental setups, we focus on the three categories with the largest number of samples: airplanes, chairs, and cars. Each shape is globally normalized and sampled to $2048$ points. The training set comprises $2832$ airplanes, $4612$ chairs, and $2458$ cars, while the evaluation set contains $405$ airplanes, $662$ chairs, and $352$ cars. Experiments are conducted on two data splits inherited from LION [LION] and TIGER [Tiger].
Metrics.
In line with the benchmarks established by PVD [PVD] and LION [LION], we adopt the 1-NN (1-nearest neighbor) accuracy [1-NN] as our evaluation criterion. This metric has demonstrated its effectiveness in assessing both the quality and diversity of the generated point clouds, with a score approaching $50\%$ indicating exceptional performance [PointFlow]. To calculate the 1-NN distance matrix, we employ two widely recognized metrics for measuring the distance between point clouds: Chamfer Distance (CD) and Earth Movers’ Distance (EMD). We also present the mean CD and the mean EMD by calculating the average CD and EMD across three distinct categories. To assess efficiency, we evaluate the training time by documenting the GPU hours and gauge the sampling efficiency by reporting the average inference time over $10$ randomly generated samples.
Figure 5: Our generation results (right) compared to baseline models (left). PointNSP generates high-quality and diverse 3D point clouds.
4.2 Performance Comparison for Single-Class Generation
Results.
To evaluate the performance of PointNSP, we compare it against several strong baseline models. Specifically, we include state-of-the-art (SoTA) diffusion-based methods: DPM [DPM], PVD [PVD], LION [LION], DIT-3D [DIT-3D], and TIGER [Tiger]. We also evaluate PVD-DDIM [DDIM], which is designed to accelerate the sampling process of PVD. In addition, we include three leading autoregressive models: PointGrow [PointGrow], CanonicalVAE [pointvqvae], and PointGPT [PointGPT]. We report results for two variants of our model—PointNSP-s and PointNSP-m—representing small and medium parameter configurations, respectively. Note that we only compare against models with publicly available code. Quality. As shown in Table 1, PointNSP consistently outperforms all baseline models on both the conventional and LION data splits, achieving state-of-the-art generation quality in terms of both CD and EMD metrics. Notably, even the lightweight PointNSP-s demonstrates competitive performance. Efficiency. Table 2 presents the training time, sampling time, and number of parameters for several strong baseline models from Table 1, including PVD, LION, TIGER, and PointGPT. In terms of training efficiency, PointNSP-s achieves the shortest training time among all selected methods, while PointNSP-m demonstrates comparable training costs to state-of-the-art diffusion-based approaches. For inference efficiency, both PointNSP-s and PointNSP-m outperform selected strong baseline methods with significantly faster sampling speeds. This advantage is primarily due to PointNSP’s parallel token generation within each scale. Additionally, PointNSP exhibits the highest parameter efficiency among all compared approaches. Notably, PointNSP-s has a smaller parameter size than PointGrow, despite the latter performing significantly worse in generation quality. Compared to diffusion-based methods, PointNSP-m offers a substantial advantage in parameter efficiency, requiring nearly half the number of parameters while maintaining superior performance.
4.3 Ablation and Analysis
Table 3: Ablation studies on the Transformer backbones, upsampling networks, token embedding, and $\operatorname{FPS}$ stochasticity.
Mask Upsampling FPS Stochasticity TE Mean CD $\downarrow$ Mean EMD $\downarrow$
Voxel PU-Net
SE 64.25 60.53
SE+A-PE 62.19 58.23
SE+A-PE 63.05 58.47
L-PE 63.22 59.71
A-PE 62.12 60.02
A-PE 61.28 57.32
SE+L-PE 60.62 57.34
[…]
Airplane Chair Car
PVD 97.53 99.88 88.37 96.37 89.77 94.89
TIGER 83.54 81.55 57.34 61.45 65.79 57.24
PointGPT 94.94 91.73 71.83 79.00 89.35 87.22
LION 86.30 77.04 66.50 63.85 64.52 54.21
PointNSP-s (ours) 78.95 68.84 58.79 55.10 59.97 52.89
[…]
Neural Architectures & Training Strategies
We conduct comprehensive ablation studies to assess the impact of various architectural components and training strategies. First, we compare two upsampling strategies: voxel-based representations and PU-Net. Our experiments show that PU-Net consistently outperforms voxel-based upsampling, owing to its permutation-equivariant design. Second, we evaluate the effectiveness of the position-aware masking strategy, which significantly boosts model performance. Third, we analyze the contribution of each embedding module in the token embedding (TE) layer, with results highlighting the notable impact of the scale embedding. We also compare learnable positional encoding (L-PE) with absolute PE (A-PE), finding that absolute PE yields better performance. Finally, we investigate the effect of $\operatorname{FPS}$ stochasticity, arising from the inherent randomness of the $\operatorname{FPS}$ algorithm. We compare against a variant with a fixed starting point, which produces deterministic downsampling. The results demonstrate that FPS stochasticity consistently enhances model generalization. These findings are summarized in Table 4, with additional ablation results—such as the effect of varying the number of scales—provided in the Appendix C.
4.4 Multi-class Generation & Dense Point Cloud Generation
Beyond the single-class generation setting, we further evaluate all models on the more challenging multi-class generation task introduced by LION [LION], aiming to assess their ability to generalize across diverse object categories. Specifically, we train PointNSP and all baseline models using class conditioning over $55$ distinct categories from ShapeNet. This conditional setup allows the model to capture class-specific characteristics while also leveraging shared geometric patterns across categories. As shown in Table 4, PointNSP significantly outperforms all strong baselines. We provide visualizations of diverse shape classes in Appendix E. We further evaluate all models on high-resolution point cloud generation with up to $8192$ points. Empirical results demonstrate that PointNSP maintains superior performance across all major metrics, with a widening performance margin as resolution increases—particularly training and inference speed (Results are shown in Table 5 (right) and the Appendix D).
4.5 Point Cloud Completion & Upsampling
We evaluate PointNSP on two key downstream tasks: point cloud completion and upsampling. For completion, we follow the experimental setup of PVD [PVD]. For upsampling, we use a factor of $r=2$, increasing the input point clouds from $1024$ to $2048$ points. As shown in Table 5 (Left &\& Middle), PointNSP consistently achieves the best performance on point cloud completion. For the upsampling task, it outperforms all selected baselines across both metrics, highlighting its effectiveness on diverse downstream applications and its potential as a foundation model. Visualization results for these two tasks are presented in Figure 7 and Appendix E.
Table 5: (Left) The shape completion task in PVD; (Middle) Our own upsampling task by upsampling to denser point clouds; (Right) Denser point cloud generation with $8192$ points.
  Category Model CD $\downarrow$ EMD $\downarrow$
Airplane SoftFlow 40.42 11.98
PointFlow 40.30 11.80
DPF-Net 52.79 11.05
PVD 44.15 10.30
[…]
Chair SoftFlow 27.86 32.95
PointFlow 27.07 36.49
DPF-Net 27.63 33.20
PVD 32.11 29.39
[…]
Car SoftFlow 18.50 27.89
PointFlow 18.03 28.51
DPF-Net 13.96 23.18
PVD 17.74 21.46
[…]
  Category Model CD $\downarrow$ EMD $\downarrow$
Airplane PVD 73.56 71.65
TIGER 71.65 59.94
PointGPT 72.11 60.12
LION 70.41 59.65
[…]
Chair PVD 53.81 64.61
TIGER 52.80 52.98
PointGPT 53.75 53.21
LION 53.98 54.33
[…]
Car PVD 58.95 48.43
PointGPT 57.26 47.85
TIGER 57.90 48.01
LION 57.14 47.56
[…]
  Category Model CD $\downarrow$ EMD $\downarrow$
Airplane PVD 69.77 69.98
TIGER 68.48 60.24
PointGPT 69.29 61.56
LION 68.95 60.38
[…]
Chair PVD 52.56 51.33
TIGER 51.87 51.85
PointGPT 52.43 52.09
LION 53.76 53.45
[…]
Car PVD 54.19 46.55
PointGPT 54.97 45.20
TIGER 53.78 44.12
LION 54.98 44.67
[…]
5 Conclusions
In this study, we present PointNSP, a novel autoregressive framework for high-quality 3D point cloud generation. Unlike prior methods, PointNSP adopts a coarse-to-fine strategy that captures the multi-scale LoD of 3D shapes without reducing them to sequences of local predictions, thereby preserving global structural coherence throughout generation. Our approach consistently outperforms existing point cloud generative models in quality while delivering significant improvements in parameter efficiency, training efficiency, and inference speed. Looking ahead, we plan to scale PointNSP toward foundation-level models on large datasets and explore its integration with diffusion-based approaches as a promising future direction.
Appendix
Appendix A Permutation Invariance of the Probability Distribution
Here we show why the distribution $p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K})=\prod_{k=1}^{K}p(\mathbf{X}_{k}|\mathbf{X}_{k-1},\dots,\mathbf{X}_{2},\mathbf{X}_{1})$ in Eq 3 preserves the permutation-invariance $p(\pi(\mathbf{x}_{1},\dots,\mathbf{x}_{N}))=p(\mathbf{x}_{1},\dots,\mathbf{x}_{N}),\forall\ \pi\in S_{N}$ in Eq 2.
By definition, the joint distribution factorizes as:
$p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K})=p(\mathbf{X}_{1})p(\mathbf{X}_{2}|\mathbf{X}_{1})\dots p(\mathbf{X}_{K}|\mathbf{X}_{1},\dots\mathbf{X}_{K-1}).$ (10)
Suppose the feature encoder and the conditional model for each scale are permutation-equivariant or permutation-invariant. Then, for any permutation $\pi_{k}$ of points within $\mathbf{X}_{k}$:
$p(\pi_{k}(\mathbf{X}_{k})|\mathbf{X}_{1},\dots\mathbf{X}_{k-1})=p(\mathbf{X}_{k}|\mathbf{X}_{1},\dots\mathbf{X}_{k-1}).$ (11)
This holds for each $k=1,\dots,K$.
$\pi=(\pi_{1},\pi_{2},\dots,\pi_{K}),\qquad\pi_{k}\in S_{s_{k}}.$ (12)
Then the joint distribution under this permutation is
$p(\pi(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K}))$ $=\prod_{k=1}^{K}p(\pi_{k}(\mathbf{X}_{k})|\mathbf{X}_{k-1},\dots,\mathbf{X}_{2},\mathbf{X}_{1})$ (13)
$=\prod_{k=1}^{K}p(\mathbf{X}_{k}|\mathbf{X}_{k-1},\dots,\mathbf{X}_{2},\mathbf{X}_{1})\ \ \text{(by permutation invariance at each scale)}$
$=p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K})$
The core operations in PointNSP —such as the feature encoder, $\operatorname{upsampling}$, and $\operatorname{FPS}$—are all designed to be permutation-equivariant or permutation-invariant. As a result, the autoregressive factorization preserves permutation invariance: permuting the points within any scale does not alter the joint probability.
Appendix B Algorithm Details
Coordinate-based Positional Encoding.
[…]
$p$ $=\lambda^{2}*z_{i}+\lambda*y_{i}+x_{i}$ (15)
$\mathbf{P}_{k}(p,2i)=\sin(\frac{p}{10000^{\frac{2i}{D}}})$ (16)
$\mathbf{P}_{k}(p,2i+1)=\cos(\frac{p}{10000^{\frac{2i}{D}}}),$ (17)
[…]
Intra-Scale Token Embedding Layer.
In this way, we derive the absolution positional encoding $\mathbf{p}^{i}_{k}$ for each token $q_{k}^{i}$ with predicted $\mathbf{\hat{X}_{k}}$. Additionally, the model needs to know which scale that each token belongs to. Therefore, $\mathbf{s}_{k}$ is a simple one-hot encoding $\mathbf{s}_{k}=\operatorname{one-hot}(k)$ out of total $K$ scales. Tokens ($q_{k}^{i}$, $q_{k}^{j}$) from the same scale $k$ share the same scale embedding $\mathbf{s}_{k}$ ($\mathbf{s}^{i}_{k}=\mathbf{s}^{j}_{k}$). Following the implementation of Llama 3 [Llama3], we add both the positional embedding $\mathbf{p}^{i}_{k}$ and the scale embedding $\mathbf{s}^{i}_{k}$ to the query $\mathbf{u}_{k}^{i}$ and key vectors $\mathbf{v}_{k}^{i}$ derived in the attention mechanism. Then the token embedding operation for each token $q^{i}_{k}$ is as follows:
where $\mathbf{W}_{\mathbf{U}}$ and $\mathbf{W}_{\mathbf{V}}$ are projection matrices for queries and keys respectively. Then the queries and keys will be upsampled to $\mathbf{U}_{k},\mathbf{V}_{k}=\operatorname{upsampling}(\mathbf{U}_{k},\mathbf{V}_{k})\in\mathbb{R}^{s_{k+1}\times d}$ following the Eq. 6, which will be fed to the transformer for next-scale token prediction.
Appendix C Hyperparameters and Reproducibility Settings
Hyper-parameters and Reproducibility Settings.
We mainly follow the setting in [LION]. Specifically, we set the learning rate $3e^{-4}$ and the batch-size 32. We perform all the experiments on a workstation with Intel Xeon Gold 6154 CPU (3.00GHz) and 8 NVIDIA Tesla V100 (32GB) GPUs.
We use an AdamW optimizer with an initial learning rate of $10^{-4}$ for VQVAE training and $10^{-3}$ for autoregressive transformer training respectively. For up-sampling and completion experiments, we follow the experimental settings of PVD [PVD].
Hyperparameter Value
# PVCNN layers 4
# PVCNN hidden dimension 1024
# PVCNN voxel grid size 32
# MLP layers 6
# Attention dimension 1024
# Attention head 32
Optimizer AdamW
Weight Decay 0.01
LR Schedule Cosine
The effect of # scales $K$.
Appendix D More Results on Dense Point Cloud Generation with 8192 points
We show sampling time and training hours of different methods in Table S1. We could clearly see that PointNSP significantly outperforms other strong methods in terms of learning efficiency and inference efficiency, which shows its strong scalability potential.
Table S1: Training time (in GPU hours), sampling time (in seconds) on 8192 points. The reported training time is averaged across three categories: airplane, chair, and car. The sampling time is averaged over $10$ random generations.
  Quality Rank Model
Training Time
(GPU hours) $\downarrow$
Sampling Time
(s) $\downarrow$
8 PointGrow 240 9.90
7 CanonicalVAE 205 8.95
6 PointGPT 296 10.56
5 PVD 201 58.1
4 LION 610 59.5
3 TIGER 320 42.1
[…]
1 PointNSP-m (ours) 190 5.48
Appendix E More Visualization Results
3D Point Clouds Generated by PointNSP.
We showcase diverse 3D point clouds generated by PointNSP, covering a wide range of shapes (Figures S2 and S3). Additional results for single-class generation across five categories are presented in Figures S4S8. We further illustrate the multi-scale sequential generation process in Figures S9S11, and provide visualizations of point cloud completion and upsampling in Figure S12.
@@ Appendix E More Visualization Results
Appendix F Other Related Works
@@
Point cloud upsampling is a crucial process in 3D modeling, aimed at increasing the resolution of low-resolution 3D point clouds. PU-Net [PU-net] pioneered the use of deep neural networks for this task, laying the foundation for subsequent advancements. Models such as PU-GCN [PU-GCN] and PU-Transformer [PU-transformer] have further refined point cloud feature extraction by leveraging graph convolutional networks and transformer networks, respectively. Additionally, approaches like Dis-PU [DisPU], PU-EVA [PU-EVA], and MPU [MPU] have enhanced the PU-Net pipeline by incorporating cascading refinement architectures. Other methods, such as PUGeo-Net [PUGeoNet], NePs [NeuralPoints], and MAFU [MAFU], employ local geometry projections into 2D space to model the underlying 3D surface. More recent approaches have reframed up-sampling as a generation task. For instance, PU-GAN [PU-GAN] and PUFA-GAN [PUFA-GAN] leverage generative adversarial networks (GANs) to produce high-resolution point clouds. Grad-PU [Grad-PU] first generates coarse dense point clouds through nearest-point interpolation and then refines them iteratively using diffusion models. In contrast, PUDM [PUDM] directly utilizes conditional diffusion models, treating sparse point clouds as input conditions for generating denser outputs. In this work, our generative model, PointNSP, incorporates upsampling networks in both training stages, making it well-suited for enhancing downstream point cloud up-sampling tasks.
Autoregressive Generative Models.
Autoregressive modeling has demonstrated exceptional success in natural language processing, leading to the development of powerful large language models (LLMs) such as [seq2seq, GPT-4o, GPT-4, GPT-3]. Building on the success of LLMs, researchers have explored autoregressive approaches across a wide range of domains, including image generation [VQVAE, VQGAN, AR-WVQ, VAR, RQtransformer, MaskGIT], graph generation [GraphRNN], video synthesis [Video-AR], molecule generation [GraphAF, moleculartransformer], and protein sequence modeling [progen, ESMfold]. At its core, autoregressive modeling involves sequentially predicting future elements based on past outputs, making it particularly effective for audio generation tasks, where such temporal dependencies are naturally present.
Appendix G Limitations & Broader Impact
PointNSP does not present major limitations, though the main challenge lies in learning high-quality multi-scale codebook embeddings for 3D point cloud representations—a process that remains resource-intensive. Several promising directions emerge for future work. One is scaling generation to produce ultra-dense point clouds (e.g., 1010k–100100k points), which could then be converted into high-quality meshes. Another is enabling fine-grained control over local geometric details, an essential step toward practical applications. While this work does not pose immediate societal risks, potential misuse for generating harmful 3D content warrants vigilance from the broader community. From a research perspective, PointNSP makes a meaningful contribution to both generative modeling and 3D point cloud research.
@@ Appendix G Limitations & Broader Impact
Figure S12: (Left) Visualizations of point cloud completion results. (Right) Visualizations of point cloud upsampling results.

### v3

@@ 9 More Experimental Results
Table S3: The Performance (1-NNA) is evaluated based on single-class generation. The second block specifies the types of generative models used in each study. The best performance is highlighted in bold, while the second-best performance is underlined. Performance is reported on two dataset splits: the top corresponds to the random split, and the bottom corresponds to the LION split.
@@
10 More Visualization Results
We showcase diverse 3D point clouds generated by PointNSP across a wide variety of shapes (Figures S3 and S4). Additional single-class generation results for five categories are provided in Figures S5S9. We further illustrate the multi-scale sequential generation process in Figures S10S11, and present more examples of point cloud completion and upsampling in Figures S13 and S12.
@@ 10 More Visualization Results
11 Other Related Works
@@
Point cloud upsampling is a crucial process in 3D modeling, aimed at increasing the resolution of low-resolution 3D point clouds. PU-Net [PU-net] pioneered the use of deep neural networks for this task, laying the foundation for subsequent advancements. Models such as PU-GCN [PU-GCN] and PU-Transformer [PU-transformer] have further refined point cloud feature extraction by leveraging graph convolutional networks and transformer networks, respectively. Additionally, approaches like Dis-PU [DisPU], PU-EVA [PU-EVA], and MPU [MPU] have enhanced the PU-Net pipeline by incorporating cascading refinement architectures. Other methods, such as PUGeo-Net [PUGeoNet], NePs [NeuralPoints], and MAFU [MAFU], employ local geometry projections into 2D space to model the underlying 3D surface. More recent approaches have reframed upsampling as a generation task. For instance, PU-GAN [PU-GAN] and PUFA-GAN [PUFA-GAN] leverage generative adversarial networks (GANs) to produce high-resolution point clouds. Grad-PU [Grad-PU] first generates coarse dense point clouds through nearest-point interpolation and then refines them iteratively using diffusion models. In contrast, PUDM [PUDM] directly utilizes conditional diffusion models, treating sparse point clouds as input conditions for generating denser outputs. In this work, our generative model, PointNSP, incorporates upsampling networks in both two training stages, making it well-suited for enhancing downstream point cloud upsampling tasks.
12 Limitations & Broader Impact
PointNSP does not exhibit major limitations, though a primary challenge lies in learning high-quality multi-scale codebook embeddings for 3D point cloud representations, particularly in avoiding codebook collapse. Several promising directions arise for future work. One avenue is scaling generation toward ultra-dense point clouds (e.g., 1010k–100100k points), which could subsequently be converted into high-fidelity meshes. Another is enabling fine-grained control over local geometric structures, a capability crucial for practical deployment. Although this work does not present immediate societal risks, potential misuse for generating harmful 3D content warrants careful consideration by the broader community. From a research standpoint, PointNSP represents a significant contribution to both the generative modeling and 3D point cloud communities.
@@ 12 Limitations & Broader Impact
Figure S12: Visualizations of more point cloud completion results.
Figure S13: Visualizations of more point cloud upsampling results.